OpenStudy (anonymous):

Solve 2(2-y)<=-3y-2 for y. Show all work.

OpenStudy (anonymous):

\[2(2-y)\leq-3y-2\] \[4-2y\leq-3y-2\] \[-2y+3y\leq-2-4\] \[y\leq-6\]